# Learning Actionable World Models for Industrial Process Control

### v1

@@ Abstract
To go from (passive) process monitoring to active process control, an effective AI system must learn about the behavior of the complex system from very limited training data, forming an ad-hoc “digital twin” with respect to process in- and outputs that captures the consequences of actions on the process’s “world”. We propose a novel methodology based on learning world models that disentangles process parameters in the learned latent representation, allowing for fine-grained control. Representation learning is driven by the latent factors that influence the processes through contrastive learning within a joint embedding predictive architecture. This makes changes in representations predictable from changes in inputs and vice versa, facilitating interpretability of key factors responsible for process variations, paving the way for effective control actions to keep the process within operational bounds. The effectiveness of our method is validated on the example of plastic injection molding, demonstrating practical relevance in proposing specific control actions for a notoriously unstable process.
Keywords JEPA  $\cdot$ contrastive learning  $\cdot$ disentangled representations
1 Introduction
@@
Figure 1: Overview of world model training and application for process control. Top: World model training using recorded process signal pairs ($\boldsymbol{x},\boldsymbol{y}$) that have been generated in the same environment (world) but with changed control parameters (actions) $\boldsymbol{a}=a_{1},...,a_{n}$. The encoder ($Enc$) generates latent representations ($\boldsymbol{z_{x}},\boldsymbol{z_{y}}$) used for predicting the actions $\hat{\boldsymbol{a}}$ through the action prediction module ($P_{a}$). The latent predictive module ($P_{z}$) guides the learning of the latent representation by incorporating actions, ensuring the representations $\boldsymbol{z}_{x}$ and $\boldsymbol{z}_{y}$ align with the process’s operational dynamics for more accurate and interpretable predictions. Detailed modeling and training is discussed in Section 3. Bottom: Application of the trained world model in real-time process control where the model predicts actions based on a reference process signal plus an observed process signal to maintain stability despite changing external conditions or influences.
@@
Automatic process monitoring plays a central role in maintaining repeatable and stable processes with high-quality products in industrial processes [1]. Commonly, the monitoring task is formulated as an anomaly detection task, as it is a priori unknown how exactly a process may deviate from normal operations [2]. To then bring it back to its intended operational step once an anomaly is detected, it is necessary to identify and remove the cause of any detected anomaly. Existing approaches focus on anomaly detection [3] and lack interpretability, making it difficult to identify the underlying root causes. A reason for this is that monitored real-world processes such as plastic injection molding or air condition flow often lack an analytical or simulation model of the relationship between measured process signals and control parameters (actions), complicating root cause analysis for control despite reliable algorithms for anomaly detection. Thus, understanding relations between process signals and actions is crucial to transition from passive monitoring to active control.
@@
Machine learning (ML) systems have shown large potential to predict process signal drifts or disturbances before they impact operations, potentially replacing conventional controller blocks by directly interpreting sensor values as process feedback [4]. To become fully effective for process control in complex, dynamic processes, world models appear as a promising solution. Having become a branch of ML research with notable progress in recent years [5, 6, 7], they are typically understood as a learned representation of the physical world, capable of predicting the consequences of various actions on the world’s state [8]. While many world models are built as generative models of specific environments (“world” referring to any relevant part of the environment) to serve as simulations for AI agent training (such as [5, 7]), we are concerned here with actionable world models that learn consequence-aware representations of world states (process signals, e.g., time-series) and directly output actions (e.g., scalars).
@@
In this paper, we introduce a novel world model for process control, inspired by joint embedding predictive architectures (JEPAs) and contrastive learning [6, 9]. Our world model makes use of recorded process signals and corresponding actions in order to learn an action-aware latent representation, facilitating disentangled action prediction in the presence of many possible control parameter adjustments. Experimental results on a real-world use case from plastic injection molding control demonstrate that our world model achieves its goals with high sample efficiency. Our main contributions are: (a) a world model for process control, based on a novel architecture that, inspired by the JEPA framework and the principle of contrastive learning, conditions the learned latent representation of the world’s state on the difference in parameters evoking this state as compared to a reference state; (b) experimental validation on a real-world injection molding case, where the contrastive modeling of pressure curve pairs and corresponding machine parameters allows the model to better distinguish between subtle variations in the data, showing disentanglement for different actions in the embedding space for the first time as compared to the state of the art; (c) a novel evaluation scheme for the quality of disentanglement using angle and distance deviations, based on the deviation between ground truth and predicted actions and showing the validity of our approach for real-world process control.
2 Background and related work
Contrastive learning. In this form of self-supervised learning (SSL) [10], a model is trained to bring embeddings of a sample and a transformed version thereof close to each other, while non-corresponding samples are pushed away from each other. Early work includes Siamese networks [11]. Contrastive methods have been introduced for time-series analysis, e.g., [12, 13], including work that disentangles variant and invariant factors in a pair of observations [14].
@@ 2 Background and related work
Generative and world modeling. Variational Auto-Encoders (VAEs) learn regularized low dimensional latent representation by using a decoder-encoder architecture. Yan et al. adapt a convolutional VAE to learn a general representation of industrial time-series data, and add additional layers to predict corresponding control (machine) parameters [3]. The model is good in detecting anomalies by identifying data drift. However, the learned representations cannot reliably disentangle the factors that affect the time-series. Our work aims to improve on this. Keurti et al. propose the homomorphism auto-encoder, equipped with a group representation acting on its latent space to learn internal representations of sensory information that are consistent with actions that modify it [15].
@@
This actionable representation represents a form of world modeling. In general, world modeling approaches can be distinguished by the kind of representations they learn. Monolithic latent representations feature a single latent vector jointly representing both states and actions [5, 16, 17, 18, 19]. Structured representations can offer higher sample efficiency, out-of-distribution efficiency and interpretability. Among such approaches are those that allow compositionality with respect to spatial [20] or temporal [21] factors. Learning structured representations for action spaces is challenging due to the combinatorial nature of actions (one can, e.g., go ‘up’ and ‘left’ at the same time). Hence, world models often fall back on monolithic action space representations [22, 23]. We take this approach as well.
@@
Joint embedding representation learning. The quality of input reconstruction from a generative model does not necessarily correspond to the quality of the learned representations [24]. JEPAs have emerged as a promising SSL approach for representation learning, offering an alternative to generative models: Instead of taking a loss in the reconstructed space, they predict and compute the loss in latent space, conditioned on actions. Recent self-supervised variants (I-JEPA [25] and V-JEPA [26]) predict masked parts of the input, learning representations for computer vision tasks. We extend the JEPA by moving away from an implicit masking action ($\boldsymbol{x}\xrightarrow{\boldsymbol{masking}}\boldsymbol{y}$) to explicitly providing arbitrary actions in a factored form ($\boldsymbol{x}\xrightarrow{\boldsymbol{a}}\boldsymbol{y}$), facilitating to learn a rich latent representation of cause and effect economically. Thus, we extend the typical SSL framework by not being agnostic to external factors (like machine parameters).
@@
Instead, our approach is action-aware self-supervised learning, combining the strengths of self-supervised feature extraction with task-specific guidance from actions (control parameters). Our world model uses them to directly guide representation learning, ensuring that the learned embeddings are actionable. This avoids the need for contrastive pairs or negative sampling strategies, reducing computational overhead and simplifying the training process. Control parameters (like e.g. machine parameters in injection molding) provide explicit domain knowledge, making the representations interpretable and fit for process control.
@@
Disentangled representation learning for time-series. Deep learning has been proposed to learn effective representations of the time-series data for downstream applications [27, 28, 12, 29]. If going to be useful for fine-grained process control, the learned representations need to be disentangled: A low-dimensional latent representation $\boldsymbol{z}=\{z_{1},z_{2},\ldots,z_{n}\}$ is sought such that all pairs of latent variables in $Z$ are independent, i.e., the change of the sequential patterns in the input corresponding to variable $z_{i}$ is invariant to variable $z_{j}$, denoted as $z_{i}\perp\!\!\!\perp z_{j}$, $\forall i,j$, where $i,j\in\{1,2,\ldots,n\}$ and $i\neq j$. The disentanglement can be extended to group disentanglement so that a group of latent variables is independent of other different groups. Most of the most of the representation learning rely on unsupervised learning due to the label scarcity. However, unsupervised learning of disentangled representations is fundamentally impossible without inductive biases on both the models and the data[30]. While interpretable semantic concepts often depend on the interaction of multiple factors rather than individual components, Li et al. argue that it is essential to interpret time-series data through a single latent representation in industrial applications [31]. A disentangled representation can be decomposed into separate dimensions, where the change in each dimension reflects a distinctive change in the real world. Li et al. propose to learn interpretable time-series representations through encouraging high mutual information between the latent codes and the original time-series to achieve disentangled latent representations [31]. In a different approach, Hamaguchi et al. [14] use contrastive learning to learn common and specific features during representation learning. Then, only the common features are used for downstream tasks. In contrast, we address disentanglement through action-conditioned world modeling.
3 Methodology
@@
Figure 2: Proposed actionable world model architecture (general case; inputs for injection molding example shown, cp. Sec. 4). There are $3$ main components in the model, encoder $Enc$, latent predictor $P_{z}$, and action predictor $P_{a}$. Losses are taken in representation space and for the action prediction.
@@
Latent predictor. The latent predictor $P_{z}$ takes as input the concatenation of the latent representation $\boldsymbol{z_{x}}$ and the transformed action $f(\boldsymbol{a})$. in our implementation, $f(\cdot)$ is simply identity function.
$P_{z}$ predicts the representation $\boldsymbol{z_{y}}$ that follows from repeating the process (that led to $\boldsymbol{x}$) after applying the change in control parameters ($\boldsymbol{a}$), modeling the transition between states caused by an action: $\boldsymbol{x}\xrightarrow{\boldsymbol{a}}\boldsymbol{y}$.
Action predictor. The action predictor $P_{a}$ uses the latent process signal representations ($\boldsymbol{z_{x}},\boldsymbol{z_{y}}$) to predict the corresponding action that led to the change. Specifically, it predicts $\boldsymbol{{\hat{a}}}=P_{a}(\boldsymbol{z_{y}}-\boldsymbol{z_{x}})$.
@@
Latent consistency loss. It enforces consistency between the predicted latent representation $\boldsymbol{\hat{z}_{y}}$ and encoded representation $\boldsymbol{z_{y}}$. This is achieved by minimizing the distance between these two representations, which is the following:
@@
Action prediction loss. It penalizes the discrepancy between the action that was predicted, $\hat{\boldsymbol{a}}$, and the action $\boldsymbol{a}$ that was actually applied. This is critical for disentangling the effect of each action dimension. The loss is defined as:
@@
4 Application to injection molding
4.1 Use case
Plastic injection molding. This is a widely utilized industrial process in which granular plastic is melted, injected into a mold, and cooled to produce components for various applications [32]. The repetitive nature of this process makes it particularly suitable for automation, but process monitoring and control is essential to ensure high product quality. Pressure sensors installed in mold cavities collect the pressure footprint of manufactured parts over production time, providing critical insight into process and product quality while enabling detection of potential anomalies [33]. Although systems exist that automatically detect such anomalies [34], identifying their root causes remains a significant challenge [3]. Root-causes are difficult to replicate in a controlled experimental setup. However, their effect on pressure signals can be mimicked by adjusting machine parameters, such as injection speed or holding pressure, which influence curve shape, scale and time shifting. For example, excess of plastic in a produced part that distorts its shape and reduces its quality, also known as flash defect, may appear as a pressure curve with higher peak than the original stable process, similar to the effect of increasing holding pressure [33]. Therefore, predicting variations in these parameters can help technicians quickly identify potential sources of anomalies and resolve them. Full automatic process control in a closed-loop fashion is however not possible in this setting for the reason that (staying in the aforementioned example) if the model predicts that pressure needs to be reduced, the correct control action is not to reduce the pressure (adjust the machine parameter), but to clean the sensor.
World modeling. We implement our methodology by adopting the deep learning architecture illustrated in Fig. 2 and detailed in Sec. 3 to an injection molding use case. The observed signals are transient pressure curves (time-series). The action space represents changes in machine parameters with respect to a reference setting (vector with a scalar per parameter). It is important to note that in this framework, the world is specific to each product being produced because the mold, the location of the sensor, and other variables impact the signal. Thus, the model must learn an ad-hoc representation of its specific world for each different production setting based on a few observations. It may rely on transfer learning from previous productions to deal with needed training data.
4.2 Datasets
Table 1: Pre-training/Fine-tuning dataset characteristics
@@ 4.2 Datasets
Data collection. We collect pressure signals from injection molding processes,following a methodology similar to [3], using Kistler cavity pressure sensors. Details of the $3$ datasets can be seen in Tab. 1.
@@
Datasets D1 and D2 differ in the specific product (plastic part) produced. Otherwise, we follow the same systematic design of experiments (DOE) where we sparsely sample signals obtained by varying a small number of machine parameters, namely holding pressure, injection speed, and mold temperature. Specifically, we systematically set each machine parameters to one of $3$ levels, resulting in a total of $273^{3}=$ 27 unique parameter combinations. We record $10$ production cycles for each machine parameter combination after reaching a steady production state, yielding in total $27\times 10=270$ samples in each of these two datasets. The $10$ cycles per parameter combination are expected to yield almost similar pressure curves during stable production conditions, hence serving as a form of data augmentation of the $27$ fundamentally different settings. An input to the model $(\mathbf{x},\mathbf{y},\mathbf{a})$ consists of a reference signal $\mathbf{x}$, an observation $\mathbf{y}$, and the action $\mathbf{a}$ (the difference between the machine parameters of the observed and the reference signal).
@@
Fig. 3 illustrates the DOE behind D1 and D2. Seeing the 3D structure of the resulting machine parameter space helps in comprehending the considerations behind dataset construction.
@@
Figure 3: Illustration of the three-dimensional machine parameter space of D1/D2 with a normalized scale. The gray points represent $27$ different machine parameters settings. The green and red points represent the most extreme settings in the corners of the cube (outline indicated by blue lines for clarity), in which individual parameters take on either minimal or maximal values. The pressure curves shown in the dashed rectangles show three examples of a reference signal $\mathbf{x}$ (green) and observed signal $\mathbf{y}$ (red) associated machine parameter setting; the arrows represent an action to transition from $\mathbf{x}\xrightarrow{\mathbf{a}}\mathbf{y}$.
@@
Dataset D3 represents data from producing a third product and includes $3$ more machine parameters in addition to the previous $3$: hot runner temperature, dosing speed, and holding time. The parameters have been varied less systematically, resembling more a collection “in the wild” as it would happen during normal operations for the sake of setting up the process. This resulted in a total of $46$ different actions, with $20$ cycles repeated per setting. By carefully analyzing the resulting pressure curves manually, we filter out $23$ of these actions that do not result in a stable process (i.e., their $20$ cycles vary significantly). Overall, this yields $23\times 20=460$ pressure curves (that are primarily explained by varying the original $3$ parameters, with the additional $3$ playing a noise-like function for our purposes).

### v2

@@ Abstract
To go from (passive) process monitoring to active process control, an effective AI system must learn about the behavior of the complex system from very limited training data, forming an ad-hoc “digital twin” with respect to process inputs and outputs that captures the consequences of actions on the process’s “world”. We propose a novel methodology based on learning world models that disentangles process parameters in the learned latent representation, allowing for fine-grained control. Representation learning is driven by the latent factors influencing the processes through contrastive learning within a joint embedding predictive architecture. This makes changes in representations predictable from changes in inputs and vice versa, facilitating interpretability of key factors responsible for process variations, paving the way for effective control actions to keep the process within operational bounds. The effectiveness of our method is validated on the example of plastic injection molding, demonstrating practical relevance in proposing specific control actions for a notoriously unstable process.
Index Terms:
JEPA, contrastive learning, disentangled representations
I Introduction
@@
Figure 1: Overview of world model training and application for process control. Top: World model training using recorded process signal pairs ($\boldsymbol{x},\boldsymbol{y}$) that have been generated in the same environment (world) but with changed control parameters (actions) $\boldsymbol{a}=a_{1},...,a_{n}$. The encoder ($Enc$) generates latent representations ($\boldsymbol{z_{x}},\boldsymbol{z_{y}}$) used for predicting the actions $\hat{\boldsymbol{a}}$ through the action prediction module ($P_{a}$). The latent predictive module ($P_{z}$) guides the learning of the latent representation by incorporating actions, ensuring the representations $\boldsymbol{z}_{x}$ and $\boldsymbol{z}_{y}$ align with the process’s operational dynamics for more accurate and interpretable predictions. Detailed modeling and training is discussed in Sec. III. Bottom: Application of the trained world model in real-time process control where the model predicts actions based on a reference process signal plus an observed process signal to maintain stability despite changing external conditions or influences.
@@
Automatic process monitoring plays a central role in maintaining repeatable and stable processes with high-quality products in industrial processes [1]. Commonly, the monitoring task is formulated as an anomaly detection task, as it is a priori unknown how exactly a process may deviate from normal operations [2]. To then bring it back to its intended operational step once an anomaly is detected, it is necessary to identify and remove the cause of any detected anomaly. Existing approaches focus on anomaly detection [3] and lack interpretability, making it difficult to identify the underlying root causes. One reason is that real-world monitored processes—such as plastic injection molding or air condition flow—often lack an analytical or simulation model of the relationship between measured process signals and control parameters (actions: physical signals, such as mold temperature adjustments, directly applied to the system, e.g., an injection molding machine, to regulate its behavior). This complicates root cause analysis for control despite reliable algorithms for anomaly detection. Thus, understanding relations between process signals and actions is crucial to transition from passive monitoring to active control.
@@
Machine learning (ML) systems have shown large potential to predict process signal drifts or disturbances before they impact operations, potentially replacing conventional controller blocks by directly interpreting sensor values as process feedback [4]. To become fully effective for process control in complex, dynamic processes, world models appear as a promising solution. Having become a branch of ML research with notable progress in recent years [5, 6, 7], they are typically referred to as a learned representation of the physical world, capable of predicting the consequences of various actions on the world’s state [8]. While many world models are built as generative models of specific environments (“world” referring to any relevant part of the environment) to serve as simulations for AI agent training (such as [5, 7]), we are concerned here with actionable world models that learn consequence-aware representations of world states (process signals, e.g., time-series data) and directly output control actions (e.g., scalars).
@@
In this paper, we introduce a novel world model for process control, inspired by joint embedding predictive architectures (JEPAs) and contrastive learning [6, 9]. Our world model makes use of recorded process signals and corresponding actions to learn an action-aware latent representation, facilitating disentangled action prediction in the presence of many possible control parameter adjustments. Experimental results on a real-world use case from plastic injection molding control demonstrate that our world model achieves its goals with high sample efficiency (limited industrial data of only 80 samples). Our main contributions are: (a) a world model for process control, based on a novel architecture, inspired by the JEPA framework and contrastive learning, conditions the learned latent representation of the world’s state on the difference in parameters evoking this state as compared to a reference state; (b) experimental validation on a real-world injection molding case, where the contrastive modeling of pressure curve pairs and corresponding machine parameters allows the model to better distinguish between subtle variations in the data, showing disentanglement for different actions in the embedding space for the first time as compared to the state of the art; (c) a novel evaluation scheme for the quality of disentanglement, based on angle and distance deviations between ground truth and predicted actions and showing the validity of our approach for real-world process control.
II Background and related work
Contrastive learning. In this form of self-supervised learning (SSL) [10], a model is trained to bring the embeddings of a sample and a transformed version thereof close to each other, while non-corresponding samples are pushed away from each other. Early work includes Siamese networks [11]. Contrastive methods have been introduced for time-series analysis, e.g., [12, 13], including work that disentangles variant and invariant factors in a pair of observations [14].
@@ II Background and related work
Generative and world modeling. Variational Auto-Encoders (VAEs) learn regularized low-dimensional latent representation by using a encoder-decoder architecture. Yan et al. adapt a convolutional VAE to learn a general representation of industrial time-series data, and add additional layers to predict corresponding control (machine) parameters [3]. The model is good in detecting anomalies by identifying data drift. However, the learned representations cannot reliably disentangle the factors that affect the time-series. Our work aims to improve on this. Keurti et al. propose the homomorphism auto-encoder, equipped with a group representation acting on its latent space to learn internal representations of sensory information that are consistent with actions that modify it [15].
@@
This actionable representation represents a form of world modeling. In general, world modeling approaches can be distinguished by the kind of representations they learn. Monolithic latent representations feature a single latent vector jointly representing both states and actions [5, 16, 17, 18, 19]. Structured representations however can offer higher sample efficiency, out-of-distribution efficiency and interpretability. Among such approaches are those that allow compositionality with respect to spatial [20] or temporal [21] factors. Learning structured representations for action spaces is challenging due to the combinatorial nature of actions (one can, e.g., go ‘up’ and ‘left’ at the same time). Hence, world models often fall back on monolithic action space representations [22, 23]. We take this approach as well.
@@
Joint embedding representation learning. The quality of input reconstruction from a generative model does not necessarily correspond to the quality of the learned representations [24]. JEPAs have emerged as a promising SSL approach for representation learning, offering an alternative to generative models: Instead of taking a loss in the reconstructed space, they predict and compute the loss in latent space, conditioned on actions. Recent self-supervised variants (I-JEPA [25] and V-JEPA [26]) predict masked parts of the input, learning representations for computer vision tasks. We extend the JEPA framework by moving away from an implicit masking action ($\boldsymbol{x}\xrightarrow{\boldsymbol{masking}}\boldsymbol{y}$) to explicitly providing arbitrary actions in a factored form ($\boldsymbol{x}\xrightarrow{\boldsymbol{a}}\boldsymbol{y}$), facilitating learning a rich latent representation of cause and effect economically. Thus, we extend the typical SSL framework by not being agnostic to external factors (like machine parameters).
@@
Instead, our approach is action-aware self-supervised learning, combining the strengths of self-supervised feature extraction with task-specific guidance from actions (control parameters). Our world model uses them to directly guide representation learning, ensuring that the learned embeddings are actionable. This avoids the need for contrastive pairs or negative sampling strategies, reducing computational overhead and simplifying the training process. Control parameters (e.g. machine parameters in injection molding) provide explicit domain knowledge, making the representations interpretable and well-suited for process control.
@@
Disentangled representation learning for time-series. Deep learning has been proposed to learn effective representations of the time-series data for downstream applications [27, 28, 12, 29]. To be useful for fine-grained process control, the learned representations must be disentangled: A low-dimensional latent representation $\boldsymbol{z}=\{z_{1},z_{2},\ldots,z_{n}\}$ is sought such that all pairs of latent variables in $Z$ are independent, i.e., the change of the sequential patterns in the input corresponding to variable $z_{i}$ is invariant to variable $z_{j}$, denoted as $z_{i}\perp\!\!\!\perp z_{j}$, $\forall i,j$, where $i,j\in\{1,2,\ldots,n\}$ and $i\neq j$. The disentanglement can be extended to group disentanglement so that a group of latent variables is independent of other different groups. Most representation learning relies on unsupervised learning due to the label scarcity. However, unsupervised learning of disentangled representations is fundamentally impossible without inductive biases on both the models and the data [30]. While interpretable semantic concepts often depend on the interaction of multiple factors rather than individual components, Li et al. argue that it is essential to interpret time-series data through a single latent representation in industrial applications [31]. A disentangled representation can be decomposed into separate dimensions, where the change in each dimension reflects a distinctive change in the real world. Li et al. propose to learn interpretable time-series representations through encouraging high mutual information between the latent codes and the original time-series to achieve disentangled latent representations [31]. Hamaguchi et al. [14] use contrastive learning to learn common and specific features during representation learning. Then, only the common features are used for downstream tasks. In contrast, we address disentanglement through action-conditioned world modeling.
III Methodology
@@
Figure 2: Proposed actionable world model architecture (general case; inputs for injection molding example shown, see Sec. IV). There are $3$ main components in the model, encoder $Enc$, latent predictor $P_{z}$, and action predictor $P_{a}$. Losses are taken in representation space and for the action prediction.
@@
Latent predictor. The latent predictor $P_{z}$ takes as input the concatenation of the latent representation $\boldsymbol{z_{x}}$ and the transformed action $f(\boldsymbol{a})$. In our implementation, $f(\cdot)$ is simply the identity function.
$P_{z}$ outputs $\boldsymbol{\hat{z}_{y}}$, the prediction of the representation $\boldsymbol{z_{y}}$ that follows from repeating the process (that led to $\boldsymbol{x}$) after applying the change in control parameters ($\boldsymbol{a}$), modeling the transition between states caused by an action: $\boldsymbol{x}\xrightarrow{\boldsymbol{a}}\boldsymbol{y}$.
Action predictor. The action predictor $P_{a}$ takes the difference between the latent representations of the process signals ($\boldsymbol{z_{y}}-\boldsymbol{z_{x}}$) as input and predicts the corresponding control action $\boldsymbol{\hat{a}}$ that induced the observed change. Formally, this is expressed as: $\boldsymbol{{\hat{a}}}=P_{a}(\boldsymbol{z_{y}}-\boldsymbol{z_{x}})$.
@@
Latent consistency loss. It enforces consistency between the predicted latent representation $\boldsymbol{\hat{z}_{y}}$ and the encoded representation $\boldsymbol{z_{y}}$. This is achieved by minimizing the distance between these two representations, which can be expressed as:
@@
Action prediction loss. It penalizes the discrepancy between the predicted action $\hat{\boldsymbol{a}}$, and the action $\boldsymbol{a}$ that is actually applied. This is critical for disentangling the effect of each action dimension. The loss is defined as:
@@
IV Application to injection molding
IV-A Use case
Plastic injection molding. This is a widely utilized industrial process in which granular plastic is melted, injected into a mold, and cooled to produce components for various applications [32]. The repetitive nature of this process makes it particularly suitable for automation, but process monitoring and control is essential to ensure high product quality. Pressure sensors installed in mold cavities collect the pressure footprint of manufactured parts over production time, providing critical insight into process and product quality while enabling detection of potential anomalies [33]. Although systems exist that automatically detect such anomalies [34], identifying their root causes remains a significant challenge [3]. Root causes are difficult to replicate in a controlled experimental setup. However, their effect on pressure signals can be mimicked by adjusting machine parameters, such as injection speed or holding pressure, which influence curve shape, scale and time shifting. For example, excess of plastic in a produced part that distorts its shape and reduces its quality, also known as flash defect, may appear as a pressure curve with higher peak than the original stable process, similar to the effect of increasing holding pressure [33]. Therefore, predicting variations in these parameters can help technicians quickly identify potential sources of anomalies and resolve them. Full automatic process control in a closed-loop fashion is however not possible in this setting for the reason that (staying in the aforementioned example) if the model predicts that pressure needs to be reduced, the correct control action is not to reduce the pressure (adjust the machine parameter), but to clean the sensor.
World modeling. We implement our methodology by adopting the deep learning architecture illustrated in Fig. 2 and detailed in Sec. III to an injection molding use case. The observed signals are transient pressure curves (time-series). The action space represents changes in machine parameters with respect to a reference setting (vector with a scalar per parameter). It is important to note that in this framework, the world is specific to each product being produced, as the mold, the location of the sensor, and other variables impact the signal. Thus, the model must learn an ad-hoc representation of its specific world for each different production setting, based on a few observations. It may rely on transfer learning from previous productions to deal with data needs economically.
IV-B Datasets
TABLE I: Pre-training/Fine-tuning dataset characteristics
@@ IV-B Datasets
Data collection. We collect pressure signals from injection molding processes, following a methodology similar to [3], using Kistler cavity pressure sensors. Details of the $3$ datasets can be seen in Tab. I.
@@
Datasets D1 and D2 differ in the specific products (plastic parts) produced. Otherwise, we follow the same systematic design of experiments (DOE), in which we sparsely sample signals obtained by varying a small number of machine parameters, namely holding pressure, injection speed, and mold temperature. Specifically, we systematically set each machine parameter to one of $3$ levels, resulting in a total of $273^{3}=$ 27 unique parameter combinations. We record $10$ production cycles for each machine parameter combination after reaching a steady production state, yielding in total $27\times 10=270$ samples in each of these two datasets. The $10$ cycles per parameter combination are expected to yield almost similar pressure curves during stable production conditions, hence serving as a form of data augmentation of the $27$ fundamentally different settings. An input to the model $(\boldsymbol{x},\boldsymbol{y},\boldsymbol{a})$ consists of a reference signal $\boldsymbol{x}$, an observation $\boldsymbol{y}$, and the action $\boldsymbol{a}$ (the difference between the machine parameters of the observed and the reference signal).
@@
Fig. 3 illustrates the DOE behind D1 and D2. Visualizing the 3D structure of the resulting machine parameter space helps in comprehending the considerations behind dataset construction.
@@
Figure 3: Illustration of the three-dimensional machine parameter space of D1/D2 with a normalized scale. The gray points represent $27$ different machine parameters settings. The green and red points represent the most extreme settings in the corners of the cube (outline indicated by blue lines for clarity), in which individual parameters take on either minimal or maximal values. The pressure curves shown in the dashed rectangles show three examples of a reference signal $\boldsymbol{x}$ (green) and observed signal $\boldsymbol{y}$ (red) associated machine parameter setting; the arrows represent an action to transition from $\boldsymbol{x}\xrightarrow{\boldsymbol{a}}\boldsymbol{y}$.
@@
Dataset D3 represents data from the production of a third product and includes $3$ more machine parameters in addition to the previous $3$: hot runner temperature, dosing speed, and holding time. The parameters vary less systematically, resembling more a collection “in the wild” as it would happen during normal operations.
This results in a total of $46$ different actions, with $20$ cycles repeated per setting. By carefully analyzing the resulting pressure curves manually, we filter out $23$ of these actions that do not result in a stable process (i.e., their $20$ cycles vary significantly). Overall, this yields $23\times 20=460$ pressure curves (that are primarily explained by varying the original $3$ parameters, with the additional $3$ playing a noise-like function for our purposes).